# Black Hole Survival Guide: Searching for Stars in the Galactic Center That Endure Partial Tidal Disruption

### v1

@@ 2.2 Mapping into MESA
As described in Section 2.1, the remnants include solar-mass stars disrupted at 4.8 Gyr with varying values of $\beta=1$, 1.5, and 2, along with an initially $M_{i}=3\,M_{\odot}$ star disrupted at 0.3 Gyr with $\beta=4$ resulting in a $M_{f}\approx 1\,M_{\odot}$ remnant. For each remnant, we use the snapshot from the end of the FLASH simulation ($\approx 80\,t_{\rm dyn}$ from pericenter, where $t_{\rm dyn}$ is the star’s dynamical time) to map into MESA (J. A. P. Law-Smith et al., 2020a). We then use MESA’s relaxation module to create an initial MESA model with a structure that closely matches the composition and entropy of the TDE remnant from the end of the FLASH simulation. We initiate our long-term stellar evolution simulations from this model. Our relaxation process allows us to observe the TDE remnant’s further evolution from a much younger stage than is possible with other approaches to the mapping that begin following the remnant’s evolution from the first few million years onward. We refer the reader to J. A. P. Law-Smith et al. (2020a) for further details on the stopping criteria in FLASH and S. Wu et al. (2020) for the mapping procedure in MESA.
@@
We are interested in comparing the evolution of TDE survivors with (1) their progenitors and (2) undisturbed stars with masses equivalent to the final masses of the remnants. Therefore, for reference, we also run a variety of undisturbed MESA models. We use a version of the 1M_pre_ms_to_wd test suite inlist, again at $Z=0.0142$, to run the evolutionary models of the undisturbed stars. Our complete inlists with all of our parameter adjustments can be found on Zenodo (R. C. Bush et al., 2025).
[…]
@@ 3 Results
Each remnant loses a portion of its envelope to the SMBH during the disruption. For the $M_{i}=1\,M_{\odot}$ progenitor star, a TDE results in stellar mass loss of 4.4% for $\beta=1$, 29.2% for $\beta=1.5$ , and 61.4% for $\beta=2$. For the $M_{i}=3\,M_{\odot}$ progenitor star, a TDE of $\beta=4$ results in stellar mass loss of 66.6%.
In Section 3.1, we will first examine the early rapid evolution that occurs during the initial bright phase of the TDE remnants. In Section 3.2, we detail the effects of the TDE on remnants’ colors and surface abundancse, and in Section 3.3, we describe the stable behavior that the TDE remnants settle into over long timescales.
@@ 3.1 Early Stellar Evolution
Zooming in on the early behavior for the remnants resulting from disruption of a $M_{i}=1\,M_{\odot}$ progenitor, Figure 2 shows that rapid evolution occurs in the remnants’ luminosities, effective temperatures, and radii during the first $10^{3}$ yr after disruption. The injection of energy from the tidal encounter causes the luminosities of the remnants (solid red) to soar orders of magnitude higher than that of their respective progenitors (black dotted lines): the luminosity of the $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant increases by a factor of $\approx 500$, and the luminosities of the $\beta=1.5$, $M_{i}=1\,M_{\odot}$ and $\beta=2$, $M_{i}=1\,M_{\odot}$ remnants increase by a factor of $\approx 10$.
We note that the lowest $\beta$ remnant experiences the highest jump in luminosity, which we attribute to the fact that its radius increases greatly while retaining almost all of its original mass. The lowest $\beta$ remnant also experiences the steepest decline in luminosity, which relates to where the energy was deposited and how long it takes to be released. The instantaneous luminosity of the remnant will be $\approx$ (energy deposited)/(diffusion time) for the region that is producing the luminosity, with the outer layers with short diffusion times producing the radiation first. Even though the $\beta=2$ encounter deposits more energy overall, its initial luminosity is still lower than the initial luminosity of $\beta=1$, which implies that less energy is deposited in the envelope where the diffusion time is short for $\beta=2$ than for $\beta=1$. In mild TDEs where the star just grazes the SMBH’s tidal radius (as in $\beta=1$), most of the energy is deposited in the outer envelope where it is quick to diffuse out, whereas more severe disruption events (as for $\beta=1.5$ and 2) result in energy deposited deeper in the interior where it takes longer to leak out. This may explain why the lowest $\beta$ encounter is initially the brightest, but also fades most rapidly.
The remnants also grow 44446666 times in radius compared to the progenitor stars. The inflated envelopes they develop during the TDE can also be seen in the FLASH snapshots in Figure 1. Directly after the TDE, the $\beta=1.5$, $M_{i}=1\,M_{\odot}$ and $\beta=2$, $M_{i}=1\,M_{\odot}$ remnants exhibit surface temperatures lower than that of their progenitor as a result of their diminished mass and puffy envelope (top panels of Figure 1, columns 2-4). The $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant does not follow this trend (top panel of Figure 1, column 1), having experienced a less severe TDE leading to relatively low mass loss, and instead shows a highly elevated temperature.
@@
During the first $10^{5}$ yr, these low-mass remnants appear to be masquerading as younger, more massive stars due to their high temperatures and luminosities. This is the stage when the disrupted stars look most distinctive from their undisturbed progenitors, and although exciting, it does not last. This early appearance fades within roughly Kelvin-Helmholtz timescales as the excess energy leaks out of the remnant and it returns to the luminosity level of its progenitor (see Appendix A for additional details on the early post-disruption evolution). Of the estimated 100,000 tidally disrupted stars we expect the Milky Way to harbor (see Equation 1), the number currently traversing this brief, bright stage should be on the order of unity: for a TDE rate of $\approx 10^{-5}\,\mathrm{yr}^{-1}$, we estimate that 1 remnant can be observed in its first $10^{5}$ yr post-disruption (see Equation 2).
$10^{-5}\,\mathrm{remnants/yr}\times 10^{10}\;\mathrm{yr}=10^{5}\;\textrm{remnants}$ (1)
$10^{5}\,\mathrm{yr}\times 10^{-5}\,\mathrm{remnants/yr}=1\;\textrm{remnant}$ (2)
@@
Following this period, the remnants continue contracting for $\approx 10^{7}$$10^{8}$ yr, consistent with the core’s thermal timescale, until they become dense enough to re-ignite hydrogen core fusion, ultimately converging to mimic the MESA models of their lower mass-equivalent stars. The vast majority of TDE survivors today should have already settled into their late-stage appearance, as we discuss in Section 3.3.
@@
Figure 3 compares the early luminosity evolution of each TDE remnant from the $1\,M_{\odot}$ progenitor to that of the remnant from the $3\,M_{\odot}$ progenitor. The evolution of undisturbed stars on the PMS and early MS with $M=1\,M_{\odot}$ and $3\,M_{\odot}$ are also shown for comparison; these represent the progenitors for our four TDE remnants, and the $M=1\,M_{\odot}$ model also serves as a mass-equivalent star for the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant. The $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant is less luminous throughout its post-disruption lifetime relative to its progenitor, which can to a large extent be attributed to its 66.6% reduction in mass. Although it has the same final mass as a $1\,M_{\odot}$ star, the remnant’s luminosity is initially larger than its mass-equivalent star by a factor of a few. Its luminosity decreases during the early ages of  $\approx\,10^{7}$ yr until it ignites hydrogen fusion, then its luminosity stabilizes to be $\approx 5$ times greater than that of the mass-equivalent star on the MS.
@@
Figure 3: Luminosity evolution for the first few 100 Myr of the TDE remnants. The TDE remnants disrupted at $\beta=1$, 1.5, and 2 from progenitor mass $M_{i}=1\,M_{\odot}$ from Figure 2 are shown with dashed lines and are compared here with their $1\,M_{\odot}$ progenitor (black dash-dotted). An additional TDE remnant with final mass $1\,M_{\odot}$, formed from a $\beta=4$ disruption of a $M_{i}=3\,M_{\odot}$ progenitor, is shown (solid) to remain brighter than its mass-equivalent star (black dash-dotted), but dimmer than its progenitor (black dotted), throughout the PMS + MS. On the right, a histogram of the number of G-type objects ($N_{G}$) per luminosity bin is shown (data from B. N. Sitarski 2016).
@@
The behavior of the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant can be contrasted with the behavior of the remnants disrupted from the $M_{i}=1\,M_{\odot}$ progenitor. Each of the $\beta=1$, 1.5, and 2 remnants return to much more similar luminosities, temperatures, and radii once hydrogen is ignited when compared with their mass equivalent stars (see Figure 2). In contrast, the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant retains a longer-lasting signature of its history via a significantly enhanced luminosity during hydrogen burning. Though the $\beta=4$, $M_{i}=3\,M_{\odot}$ has a final mass of $M_{f}=1\,M_{\odot}$ that is very similar to the final mass $M_{f}=0.956\,M_{\odot}$ of the $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant, the latter experienced a less disruptive encounter and does not show a similarly enduring imprint of the TDE during the hydrogen burning phase.
@@
Figure 4: Hertzsprung-Russell diagram of simulated stars on the hydrogen-burning main sequence. Data from Gaia DR3 (Gaia Collaboration et al., 2016, 2021) of the closest 50,000 stars is plotted in the background as a logarithmic histogram where grayscale shading indicates number density, to show the natural spread of Milky Way stars. The filled circles are main sequence single stars simulated in MESA, with the ZAMS masses listed in the upper right legend. The open diamonds represent the $\beta$ = 1, 1.5, and 2, $M_{i}=1\,M_{\odot}$ TDE remnants, and the open square is the $\beta=4$, $M_{i}=3\,M_{\odot}$ TDE remnant; all are shown at snapshots when they have returned to stable hydrogen burning, and the final masses of each remnant are noted in the bottom legend. All remnants appear substantially cooler and dimmer than their respective progenitors, while appearing subtly hotter and brighter than their mass-equivalent stars (or extremely hotter and brighter, in the case of the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant).
Figure 5: Surface mass fractions of selected elements $X_{i}$ in our remnants, scaled to solar metallicity abundances $X_{i,\odot}$. We show results for the elements $X_{i}$ listed in the legend. Lines end just before core He ignition. The $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant retains near-solar abundances until a dredge-up phase when it becomes a red giant. By the main sequence, the envelopes of the disrupted stars with $\beta>1$ are enriched in nitrogen (orange dashes) and depleted in carbon (red dots), both indicators of a stellar core undergoing CNO fusion. The presence of core-like CNO ratios on the stars’ surfaces is a result of convective mixing that occurs during the first $10^{5}$ yr after disruption. Enhancement of helium is also observed in each of these models. These differences may manifest as spectroscopic signatures in the remnants of partial TDEs.
@@ 3.2 Colors and Abundances
The TDE remnants eventually come to settle into a stable period of hydrogen fusion which lasts for $10^{9}$$10^{11}$ yr, which can be interpreted as them returning to the behavior of stars on the main sequence. When placed on a Hertzsprung-Russell diagram (Figure 4), survivors are offset from what would be expected. Our models appear cooler and less luminous (redder) than their corresponding progenitor, but also hotter and more luminous (bluer) than mass-equivalent stars. For TDE remnants originating from a $M_{i}=1\,M_{\odot}$ progenitor, the offset of the TDE remnants from their mass-equivalent stars is subtle. However, the offset is more extreme for the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant, which is an order of magnitude brighter and $>1000$ K hotter than its $1\,M_{\odot}$ mass-equivalent star. This results from the remnant possessing a denser and hotter core, such that it fuses hydrogen via both pp-chain and CNO fusion in almost equal measure. However, all remnants fall within the natural deviation of stars on the main sequence observed by Gaia; therefore, without an independent mass measurement, color alone is not a conclusive method for finding TDE survivors.
@@
An independent avenue for identifying survivor stars would be searching for evidence of atmospheric enrichment. Simulations show that TDEs can effectively mix the composition of the star, dislodging helium and heavier elements such as carbon and nitrogen from the core (C. S. Kochanek, 2016; M. Gallegos-Garcia et al., 2018; J. Law-Smith et al., 2019; B. Mockler et al., 2022). Subsequent to the TDE, each remnant develops convection in the envelope as it evolves, transporting some of these heavier elements to the surface. Figure 5 shows surface abundances scaled to solar up until just before core He ignition. Focusing first on the stars disrupted from a $1\,M_{\odot}$ progenitor, the least disturbed $\beta=1$ remnant exhibits convection that is confined to the envelope. Its abundances stay roughly equivalent to solar, until it experiences dredge-up as it enters the red giant branch. However, the $\beta=1.5$ and $\beta=2$ remnants develop progressively larger convective zones in the $10^{5}$ years after disruption. Figure 1 demonstrates that the material of the $\beta=2$, $M_{i}=1\,M_{\odot}$ remnant is cooler throughout than that of the lower-$\beta$ remnants. This translates into cooler temperatures for the MESA model of the remnant, and leads to the formation of a deeper convective zone. Such an effect is in line with expectations that the depth of the convective envelope increases with decreasing ZAMS mass for undisturbed stars of the same masses as the remnants. The presence of convection in these remnants promotes further mixing of elements from the bottom of the convective zone, which include products from the CNO burning which occurred in the core of the $1\,M_{\odot}$ progenitor before disruption. Thus, the degree of element enrichment increases with the $\beta$-parameter for a given initial progenitor mass, because stars that approach closer to the SMBH experience more severe disruption that produces diffuse, cooler remnants with deeper convection.

### v2

@@ 2.2 Mapping into MESA
As described in Section 2.1, the remnants include solar-mass stars disrupted at 4.8 Gyr with varying values of $\beta=1$, 1.5, and 2, along with an initially $M_{i}=3\,M_{\odot}$ star disrupted at 0.3 Gyr with $\beta=4$ resulting in a $M_{f}\approx 1\,M_{\odot}$ remnant. For each remnant, we use the snapshot from the end of the FLASH simulation to map into MESA ($\approx 80\,t_{\rm dyn}$ from pericenter, which corresponds to $\approx 38$ hr after pericenter for the $M_{i}=1\,M_{\odot}$ progenitor and $\approx 124$ hr for the $M_{i}=3\,M_{\odot}$ progenitor) (J. A. P. Law-Smith et al., 2020a). By this point in the FLASH simulation, the remnant’s central density has relaxed to a steady state. We divide the domain of the final FLASH snapshot into material that is bound and unbound to the star; bound materials (including portions of the streams) become included as part of the remnant. We then use MESA’s relaxation module to create an initial MESA model with a structure that closely matches the composition and entropy of the TDE remnant from the end of the FLASH simulation (Appendix B, B. Paxton et al., 2018). We initiate our long-term stellar evolution simulations from this model. Our relaxation process allows us to observe the TDE remnant’s further evolution from a much younger stage than is possible with other approaches to the mapping that begin following the remnant’s evolution from the first few million years onward. We refer the reader to J. A. P. Law-Smith et al. (2020a) for further details on the stopping criteria in FLASH and S. Wu et al. (2020) for the mapping procedure in MESA.
@@
We are interested in comparing the evolution of TDE survivors with (1) their progenitors and (2) stars that are equal in mass to the remnants, which we call mass-equivalent stars. Consequently, we run a variety of MESA stellar evolutionary models of single, undisturbed stars in addition to the TDE remnant models listed in Table 1. (1) To model the progenitors, we evolve undisturbed stars with zero age main sequence (ZAMS) masses equivalent to the initial masses of the remnants before the TDE occurred (i.e., $M=1\,M_{\odot}$ and $3\,M_{\odot}$). (2) To model mass-equivalent stars, we evolve undisturbed stars with ZAMS masses equivalent to the final masses of the remnants after the TDE (i.e., $M=1,0.956,0.708,{\rm and\,}0.386\,M_{\odot}$). Our complete MESA inlists with all of our parameter adjustments can be found on Zenodo (R. C. Bush et al., 2025).
@@ 3 Results
Each remnant loses a portion of its envelope to the SMBH during the disruption, ranging from just a few percent to over two thirds mass loss, depending on the disruption parameter $\beta$ (see Table 1). In Section 3.1, we will first examine the early rapid evolution that occurs during the initial bright phase of the TDE remnants; in Section 3.2, we detail the effects of the TDE on remnants’ colors and surface abundances; and in Section 3.3, we describe the stable behavior that the TDE remnants settle into over long timescales.
@@ 3.1 Early Stellar Evolution
Zooming in on the early behavior for the remnants resulting from disruption of a $M_{i}=1\,M_{\odot}$ progenitor, Figure 2 shows that rapid evolution occurs in the remnants’ luminosities, effective temperatures, and radii during the first $10^{3}$ yr after disruption. The injection of energy from the tidal encounter causes the luminosities of the remnants (solid red lines) to soar orders of magnitude higher than that of their respective progenitors (black dotted): the luminosity of the $1M\beta 1$ remnant increases by a factor of $\approx 500$, and the luminosities of the $1M\beta 1.5$ and $1M\beta 2$ remnants increase by a factor of $\approx 10$.
We note that the $1M\beta 1$ remnant experiences the highest jump in luminosity, which we attribute to the fact that its radius increases greatly while retaining almost all of its original mass. The $1M\beta 1$ remnant also experiences the steepest decline in luminosity, which relates to where the energy was deposited and how long it takes to be released. The instantaneous luminosity of the remnant will be $\approx$ (energy deposited)/(diffusion time) for the region that is producing the luminosity, with the outer layers with short diffusion times producing the radiation first. Even though the $\beta=2$ encounter deposits more energy overall, the initial luminosity of the $1M\beta 2$ remnant is still lower than the initial luminosity of the $1M\beta 1$ remnant, which implies that less energy is deposited in the envelope where the diffusion time is short for $\beta=2$ than for $\beta=1$. In mild TDEs where the star just grazes the SMBH’s tidal radius (as for $1M\beta 1$), most of the energy is deposited in the outer envelope where it is quick to diffuse out, whereas more severe disruption events (as for $1M\beta 1.5$ and $1M\beta 2$) result in energy deposited deeper in the interior where it takes longer to leak out. This may explain why the $1M\beta 1$ encounter is initially the brightest, but also fades most rapidly.
The remnants also grow 4466 times in radius compared to the progenitor stars. The inflated envelopes they develop during the TDE can also be seen in the FLASH snapshots in Figure 1. Directly after the TDE, the $1M\beta 1.5$ and $1M\beta 2$ remnants exhibit surface temperatures lower than that of their progenitor as a result of their diminished mass and puffy envelope (top panels of Figure 1, columns 2-4). The $1M\beta 1$ remnant does not follow this trend (top panel of Figure 1, column 1), having experienced a less severe TDE leading to relatively low mass loss, and instead shows a highly elevated temperature.
@@
During the first $10^{5}$ yr, these low-mass remnants appear to be masquerading as younger, more massive stars due to their high temperatures and luminosities. This is the stage when the disrupted stars look most distinctive from their undisturbed progenitors, and although exciting, it does not last. This early appearance fades within roughly Kelvin-Helmholtz timescales as the excess energy leaks out of the remnant and it returns to the luminosity level of its progenitor (see Appendix A for additional details on the early post-disruption evolution). Of the estimated one million tidally disrupted remnants we expect the Milky Way to harbor (see Equation 1), the number currently traversing this brief, bright stage should be on the order of 10: for a partial TDE rate of $\approx 10^{-4}\,\mathrm{yr}^{-1}$, we estimate that $\lessapprox 10$ remnants can be observed in their first $10^{5}$ yr post-disruption (see Equation 2).
$10^{-4}\,\mathrm{remnants/yr}\times 10^{10}\;\mathrm{yr}=10^{6}\;\textrm{remnants}$ (1)
$10^{5}\,\mathrm{yr}\times 10^{-4}\,\mathrm{remnants/yr}=10\;\textrm{remnants}$ (2)
@@
Following this initial bright period, the remnants continue contracting for $\approx 10^{7}$$10^{8}$ yr, consistent with the core’s thermal timescale, until they become dense enough to re-ignite hydrogen core fusion, ultimately converging to mimic the MESA models of their lower mass-equivalent stars. The vast majority of TDE survivors today should have already settled into their late-stage appearance, as we discuss in Section 3.3.
@@
Figure 3 compares the early luminosity evolution of each TDE remnant from the $1\,M_{\odot}$ progenitor to that of the remnant from the $3\,M_{\odot}$ progenitor. The evolution of undisturbed stars on the PMS and early MS with $M=1\,M_{\odot}$ and $3\,M_{\odot}$ are also shown for comparison; these represent the progenitors for our four TDE remnants, and the $M=1\,M_{\odot}$ model also serves as a mass-equivalent star for the $3M\beta 4$ remnant. The $3M\beta 4$ remnant is less luminous throughout its post-disruption lifetime relative to its progenitor, which can to a large extent be attributed to its 64.7% reduction in mass. Although it has roughly the same final mass as a $1\,M_{\odot}$ star, the remnant’s luminosity is initially larger than its mass-equivalent star by a factor of a few. Its luminosity decreases during the early ages of  $\approx\,10^{7}$ yr until it ignites hydrogen fusion, then its luminosity stabilizes to be $\approx 5$ times greater than that of the mass-equivalent star on the MS.
@@
Figure 3: Luminosity evolution for the first few 100 Myr of the TDE remnants. The $1M\beta 1$, $1M\beta 1.5$, and $1M\beta 2$ TDE remnants from Figure 2 are shown with dashed lines and are compared here with their $1\,M_{\odot}$ progenitor (black dash-dotted). Additionally, the $3M\beta 4$ remnant is shown (solid) to remain brighter than its mass-equivalent star (black dash-dotted), but dimmer than its progenitor (black dotted), throughout the PMS + MS. On the right, a histogram of the number of G-type objects ($N_{G}$) per luminosity bin is shown (data from A. M. Ghez et al. 2004, 2005; S. D. Hornstein et al. 2007; K. Phifer et al. 2013; B. N. Sitarski 2016; G. Witzel et al. 2014).
@@
The behavior of the $3M\beta 4$ remnant can be contrasted with the behavior of the remnants disrupted from the $M_{i}=1\,M_{\odot}$ progenitor. The $1M\beta 1$, $1M\beta 1.5$, and $1M\beta 2$ remnants return to much more similar luminosities, temperatures, and radii once hydrogen is ignited when compared with their mass-equivalent stars (see Figure 2). In contrast, the $3M\beta 4$ remnant retains a longer-lasting signature of its history via a significantly enhanced luminosity during hydrogen burning. Though the $3M\beta 4$ remnant has a final mass of $M_{f}=1.06\,M_{\odot}$ that is very similar to the final mass $M_{f}=0.956\,M_{\odot}$ of the $1M\beta 1$ remnant, the latter experienced a less disruptive encounter and does not show a similarly enduring imprint of the TDE during the hydrogen burning phase.
@@
Figure 4: Hertzsprung-Russell diagram of simulated stars on the hydrogen-burning main sequence. Data from Gaia DR3 (Gaia Collaboration et al., 2016, 2021) of the closest 50,000 stars is plotted in the background as a logarithmic histogram where grayscale shading indicates number density, to show the natural spread of Milky Way stars. The filled circles are main sequence single stars simulated in MESA, with the ZAMS masses listed in the upper right legend. The open diamonds represent the $1M\beta 1$, $1M\beta 1.5$, and $1M\beta 2$ TDE remnants, and the open square is the $3M\beta 4$ TDE remnant; all are shown at snapshots when they have returned to stable hydrogen burning, and the final masses of each remnant are noted in the bottom legend. All remnants appear substantially cooler and dimmer than their respective progenitors, while appearing subtly hotter and brighter than their mass-equivalent stars (or extremely hotter and brighter, in the case of the $3M\beta 4$ remnant).
Figure 5: Surface mass fractions of selected elements $X_{i}$ in our remnants, scaled to solar metallicity abundances $X_{i,\odot}$. We show results for the elements $X_{i}$ listed in the legend. Lines end just before core He ignition. The $1M\beta 1$ remnant retains near-solar abundances until a dredge-up phase when it becomes a red giant. For the disrupted stars with $\beta>1$, by the main sequence the envelopes are enriched in nitrogen (orange dashes) and depleted in carbon (red dots), both indicators of CNO fusion processes in the core. The presence of core-like CNO ratios on the stars’ surfaces is a result of convective mixing that occurs during the first $10^{5}$ yr after disruption. Enhancement of helium is also observed in each of these models. These differences may manifest as spectroscopic signatures in the remnants of partial TDEs.
@@ 3.2 Colors and Abundances
The TDE remnants eventually come to settle into a stable period of hydrogen fusion which lasts for $10^{9}$$10^{11}$ yr, which can be interpreted as them returning to the behavior of stars on the main sequence. When placed on a Hertzsprung-Russell diagram (Figure 4), survivors are offset from what would be expected. Our models appear cooler and less luminous (redder) than their corresponding progenitor, but also hotter and more luminous (bluer) than mass-equivalent stars. For TDE remnants originating from a $M_{i}=1\,M_{\odot}$ progenitor, the offset of the TDE remnants from their mass-equivalent stars is subtle. However, the offset is more extreme for the $3M\beta 4$ remnant, which is an order of magnitude brighter and >1000>1000 K hotter than its $1\,M_{\odot}$ mass-equivalent star. This results from the remnant possessing a denser and hotter core, such that it fuses hydrogen via both pp-chain and CNO fusion in almost equal measure. However, all remnants fall within the natural deviation of stars on the main sequence observed by Gaia; therefore, without an independent mass measurement, color alone is not a conclusive method for finding TDE survivors.
@@
An independent avenue for identifying survivor stars would be searching for evidence of atmospheric enrichment. Simulations show that TDEs can effectively mix the composition of the star, dislodging helium and heavier elements such as carbon and nitrogen from the core (C. S. Kochanek, 2016; M. Gallegos-Garcia et al., 2018; J. Law-Smith et al., 2019; B. Mockler et al., 2022). Subsequent to the TDE, each remnant develops convection in the envelope as it evolves, transporting some of these heavier elements to the surface. Figure 5 shows surface abundances scaled to solar up until just before core He ignition. Focusing first on the stars disrupted from a $1\,M_{\odot}$ progenitor, the least disturbed remnant ($1M\beta 1$) exhibits convection that is confined to the envelope. Its abundances stay roughly equivalent to solar, until it experiences dredge-up as it enters the red giant branch. However, the $1M\beta 1.5$ and $1M\beta 2$ remnants develop progressively larger convective zones in the $10^{5}$ years after disruption. Figure 1 demonstrates that the material of the $1M\beta 2$ remnant in FLASH is cooler throughout than that of the lower-$\beta$ remnants. This translates into cooler temperatures for the MESA model of the remnant, and leads to the formation of a deeper convective zone. Such an effect is in line with expectations that the depth of the convective envelope increases with decreasing ZAMS mass for undisturbed stars of the same masses as the remnants. The presence of convection in these remnants promotes further mixing of elements from the bottom of the convective zone, which include products from the CNO burning which occurred in the core of the $1\,M_{\odot}$ progenitor before disruption. Thus, the degree of element enrichment increases with the $\beta$-parameter for a given initial progenitor mass, because stars that approach closer to the SMBH experience more severe disruption that produces diffuse, cooler remnants with deeper convection.
We find that this results in an envelope abnormally enriched in helium for remnants with deep convective envelopes (Figure 5) as compared to the surface helium abundance expected of typical single stars at solar metallicity. After disruption and once the remnant has reignited core hydrogen-burning, we find that the $1M\beta 1.5$ and $1M\beta 2$ remnants have enhanced He abundance relative to solar (using the M. Asplund et al. 2009 solar abundances) ranging from $X_{\rm He}/X_{\mathrm{He},\odot}\approx 1.2$$1.3$. Similarly, the envelope may become enriched in nitrogen and depleted in carbon due to mixing. Figure 5 shows that the difference in carbon and nitrogen mass fractions compared to a Sun-like star can be significant, with $X_{\rm N}/X_{\mathrm{N},\odot}\approx 3.2$ and $X_{\rm C}/X_{\mathrm{C},\odot}\approx 0.4$. This difference in envelope composition will likely affect the appearance of stellar spectra and potentially provides a secondary diagnostic tool to find these remnants. The mixing seen in MESA which manifests as each jump in the abundance in Figure 5 is sensitive to choices of convection parameters and other numerical considerations. It takes approximately $10^{5}$ yr for the mixing to conclude, after which the surface ratios remain stable until the end of the main sequence. The exact time evolution of these abundance ratio values before the mixing concludes should be taken as approximate.
The altered composition of the remnant is especially noticeable for the $3M\beta 4$ remnant, shown in the right-most panel of Figure 5. Compared to a solar-mass star at solar abundance, which has a similar total mass to this remnant, each of the hydrogen, helium, carbon, and nitrogen abundances at the surface are found to be appreciably altered ($X_{\rm He}/X_{\mathrm{He},\odot}\approx 2$, $X_{\rm N}/X_{\mathrm{N},\odot}\approx 7$, $X_{\rm C}/X_{\mathrm{C},\odot}\approx 0.2$). This is still caused by convective mixing after the TDE, but the effect is enhanced since the $3\,M_{\odot}$ progenitor primarily burned hydrogen on the main sequence via the CNO cycle before the TDE occurred. For these highly disrupted, higher-mass main sequence stars, we expect an even greater prospect of distinguishing a history of tidal disruption through an unusual stellar spectrum. Our findings are corroborated by a similar study completed by M. Sharma et al. (2024), in which they simulate the evolution of TDE remnants of $M_{i}=1$ and $M_{i}=3\,M_{\odot}$ MAMS stars and also find that high-$\beta$ remnants exhibit compositional mixing that leads to heightened surface enrichment of N and He (Figures 13, 14, and 18 in M. Sharma et al. 2024).
3.3 Long-Term Stellar Evolution
Our three remnants disrupted from a $1\,M_{\odot}$ progenitor lose varying degrees of mass during the tidal disruption event and consequently spend the rest of their lives burning less intensely. This lowered rate of fusion resulting from their mass loss allows them to live billions of years longer than their progenitor. While the $1\,M_{\odot}$ progenitor ages off of the main sequence around age 11 Gyr (roughly 6 Gyr after disruption), the $1M\beta 1$ remnant’s life is extended by an additional 3 Gyr; by 19 Gyr for the $1M\beta 1.5$ remnant; and by 140 Gyr for the $1M\beta 2$ remnant, as seen in Figure 2. Conversely, each remnant has a shorter lifetime than its mass-equivalent star, which is consistent with the remnant having the core density of a higher-mass star.
Once the $3M\beta 4$ remnant reignites core hydrogen fusion, it similarly lives longer than its progenitor and reaches the RGB earlier than its mass-equivalent star. Notably, the remnant’s luminosity during hydrogen burning is larger than that of its mass-equivalent star by a factor of $\approx 10$ during its long-term evolution on the main sequence, which is a much larger contrast than any of the remnants disrupted from a $M_{i}=1\,M_{\odot}$ progenitor achieve. Consequently, it may be possible to differentiate the TDE remnant of a heavily disrupted high-mass star (e.g., $3M\beta 4$) from the remnant of a mildly disrupted low-mass star (e.g., $1M\beta 1$), in that a more heavily disrupted high-mass remnant maintains a heightened luminosity signature for its entire life, even after having reignited core hydrogen fusion.
While our work is most concerned with the previously unstudied early-time behavior of the remnants, we evolve our models self-consistently until the end of their time on the main sequence to offer a holistic picture of their entire evolution from initial conditions to old age. We find that our early-time behavior leads naturally to late-time behavior that is consistent with M. Sharma et al. (2024), which studied the late-time evolution in detail. At late times, both of our studies find that the remnant is cooler and dimmer than its progenitor, yet hotter and brighter than its mass equivalent star; that the remnants resulting from higher-$\beta$ interactions live longer; and that the brightest remnants come from low-$\beta$ encounters that lose the least amount of mass.